# Correlation-Embedded Transformer Tracking: A Single-Branch Framework

### v1

@@ 4 Improved Single-Branch Tracking
4.1 Overview of Improvements
The overall architecture of SuperSBT is illustrated in Fig. 5. Following Principle A and Principle D, SuperSBT adopts a hierarchical transformer architecture (see Sec. 4.2) and independent layer design for shallow and deep model stages. Specifically, we stack local modeling layers (see Sec. 4.3) and patch merging layers in the first two stages, gradually reducing the number of tokens and expanding the channel dimension. It is worth noting that the initial two-stage processes involving the template and search image tokens are separated. Based on Principle B, we propose a unified relation modeling layer (see Sec. 4.4) for joint feature extraction and correlation in the third stage. Following Principle C and Principle E, we select the optimal architecture variants (see Sec. 4.9), such as layer number and channel dimension, to improve both performance and speed. Additionally, we chose patch-merging and a convolutional layer as our PaE method based on Principle F. In the final, SuperSBT is further enhanced by masked image modeling pre-training (see Sec. 4.6), a simple temporal modeling scheme (see Sec. 4.7), and an enhanced prediction head (see Sec. 4.8).
4.2 Hierarchical Architecture
@@
4.3 Local Modeling Layer
Here, we present the local modeling layer, for the shallow model stage. The structure of the local modeling layer is illustrated in Fig. 6(a). Unlike the standard transformer layer[12, 14, 15], the Multi-head attention layer is replaced by the multi-layer perceptron to enrich the local representations in the shallow model stage. The remaining part of the local modeling layer follows the standard transformer layer. The data processing of can be formulated as follows:
$\mathbf{f}_{att}&=\text{LayerNorm}(\mathbf{f}+\text{MLP}(\mathbf{f})),\quad\mathbf{f}\in\{z,x\}\\ \mathbf{f}_{out}&=\text{LayerNorm}(\mathbf{f}_{att}+\text{MLP}(\mathbf{f}_{att})),\quad\mathbf{f}\in\{z,x\}\end{aligned}$ (10)
@@ 4.3 Local Modeling Layer
4.4 Unified Relation Modeling
@@
$\mathbf{f}_{zx}&=\text{Concat}(\mathbf{f}_{z},\mathbf{f}_{x}),\\ \mathbf{f}_{zx-att}&=\text{LayerNorm}(\mathbf{f}_{zx}+\text{MHSA}(\mathbf{f}_{zx})),\\ \mathbf{f}_{zx-out}&=\text{LayerNorm}(\mathbf{f}_{zx-att}+\text{MLP}(\mathbf{f}_{zx-att})),\end{aligned}$ (11)
@@
Based on Principle B, we do not adopt any specialized local attention (e.g., window attention), as it is harmful to do parallel computing. Thus, our SuperSBT can run exceptionally fast. Moreover, we also abandon the explicit modeling of intra-/inter-image relationship modeling. We argue that the URM method can naturally learn the layer pattern of intra-/inter-image modeling for the whole network.
4.5 Relative Position Encoding
@@ 3 $8\&16$
4.6 Masked Image Modeling Pre-training
@@
4.7 Temporal Modeling Scheme
We demonstrate our temporal modeling scheme to enhance our SuperSBT tracker further. Similar to the dynamic template mechanism in Stark [43], we introduce a dynamically updated template sampled from intermediate frames as an additional input.In the main stage of SueprSBT, images of the triplet are concatenated and then sent to the FRM layers.
Beyond the spatial information from the initial template, the dynamic template can capture the target appearance changes with time, providing additional temporal information and making SuperSBT more robust to changing scenarios. The difference between Stark [43] is that our SuperSBT can directly update the image input while Stark updates the image features extracted from ResNet [11] backbone and only the decoder part models the temporal variations. Thus, our SuperSBT learns the spatio-temporal relationship among templates and search regions more thoroughly than Stark.
4.8 Mix-MLP Prediction Head
In addition to the Convoluational-based head, we propose a Mix-MLP head, which can jointly model the dependency between the spatial and channel dimensions of the input features.
@@ 4.8 Mix-MLP Prediction Head
4.9 Network Variants
Following the guidelines from Sec. 3.4, five variants of SBT are described in Tab. II. It includes two base SBT variants and three SuperSBT variants. We place most layers in the third model stage based on Principle C. Moreover, we adjust the channel numbers and layer dimensions to maintain high speed and model performance based on Principle E.
4.10 Training Loss
@@
Method Year LaSOT [62] TrackingNet [18] GOT-10k [1]
[…]
SuperSBT-Light ours 65.8 75.3 70.6 81.4 86.2 79.3 69.4 79.4 64.1
SuperSBT-Small ours 67.5 77.1 73.1 82.7 87.3 81.2 71.6 81.4 68.3
SuperSBT-Base ours 70.0 79.8 76.1 84.0 88.4 83.2 74.4 83.9 71.3
SuperSBT-Base-384 ours 72.8 82.5 78.6 84.8 88.9 83.7 75.5 84.3 72.4
Plain-SBT ours 65.7 75.4 70.1 81.2 86.1 80.3 69.2 79.1 64.0
Hi-SBT ours 65.3 74.7 70.3 81.0 85.6 79.0 69.5 79.8 63.9
SeqTrack [63] 2023 69.9 79.7 76.3 83.3 88.3 82.2 74.7 84.7 71.8
ROMTrack [64] 2023 69.3 78.8 75.6 83.6 88.4 82.7 72.9 82.9 70.2
GRM [65] 2023 69.9 79.3 75.8 84.0 88.7 83.3 73.4 82.9 70.4
OStrack [46] 2022 69.1 78.7 75.2 83.1 87.8 82.0 71.0 80.4 68.2
SimTrack [48] 2022 69.3 78.5 - 82.3 86.5 - 68.6 78.9 62.4
Mixformer [47] 2022 69.2 78.7 74.7 83.1 88.1 81.6 70.7 80.0 67.8
SwinTrack [44] 2022 67.2 - 70.8 81.1 - 78.4 71.3 81.9 64.5
TransT [4] 2022 64.9 73.8 69.0 81.4 86.7 80.3 67.1 76.8 60.9
ToMP50 [24] 2022 67.6 - - 81.2 86.4 78.9 - - -
CSWinTT [66] 2022 66.2 75.2 70.9 81.9 86.7 79.5 69.4 78.9 65.4
UTT [67] 2022 64.6 - 67.2 79.7 - 77.0 67.2 76.3 60.5
ARDiMPsuper [55] 2021 65.3 73.2 68.0 80.5 85.6 78.3 70.1 80.0 64.2
TrDiMP [42] 2021 63.9 - 61.4 78.4 83.3 73.1 68.8 80.5 59.7
TrSiam [42] 2021 62.4 - 60.0 78.1 82.9 72.7 67.3 78.7 58.6
STMTrack [32] 2021 60.6 69.3 63.3 80.3 85.1 76.7 64.2 73.7 57.5
SiamBAN-ACM [28] 2021 57.2 - - 75.3 81.0 71.2 - - -
SiamGAT [68] 2021 53.9 63.3 53.0 - - - 62.7 74.3 48.8
DSTrpn [69] 2021 43.4 51.3 - 64.9 58.9 - - - -
SiamR-CNN [70] 2020 64.8 72.2 - 81.2 85.4 80.0 64.9 72.8 59.7
Ocean [71] 2020 56.0 65.1 56.6 - - - 61.1 72.1 47.3
KYS [72] 2020 55.4 63.3 - 74.0 80.0 68.8 63.6 75.1 51.5
DCFST [23] 2020 - - - 75.2 80.9 70.0 63.8 75.3 49.8
SiamFC++ [26] 2020 54.4 62.3 54.7 75.4 80.0 70.5 59.5 69.5 47.9
PrDiMP [73] 2020 59.8 68.8 60.8 75.8 81.6 70.4 63.4 73.8 54.3
SiamAttn [31] 2020 56.0 64.8 - 75.2 81.7 - - - -
MAML [35] 2020 52.3 - - 75.7 82.2 72.5 - - -
D3S [74] 2020 - - - 72.8 76.8 66.4 59.7 67.6 46.2
SiamCAR [27] 2020 50.7 60.0 51.0 - - - 56.9 67.0 41.5
SiamBAN [28] 2020 51.4 59.8 52.1 - - - - - -
DiMP [9] 2019 56.9 65.0 56.7 74.0 80.1 68.7 61.1 71.7 49.2
SiamPRN++ [7] 2019 49.6 56.9 49.1 73.3 80.0 69.4 51.7 61.6 32.5
ATOM [3] 2019 51.5 57.6 50.5 70.3 77.1 64.8 55.6 63.4 40.2
ECO [21] 2017 32.4 33.8 30.1 55.4 61.8 49.2 31.6 30.9 11.1
MDNet [75] 2016 39.7 46.0 37.3 60.6 70.5 56.5 29.9 30.3 9.9
SiamFC [6] 2016 33.6 42.0 33.9 57.1 66.3 53.3 34.8 35.3 9.8
5 Experiments
@@ 5.1 Implementation Details
5.2 Evaluation on TrackingNet, LaSOT, and GOT-10k
@@
5.3 Evaluation on Other Datasets
@@
5.4 Ablation Study and Analysis
Table VI: Ablations on the effects of architecture variants: network depth $N_{D}$, width $N_{W}$, and model structure (plain/hierarchical). $N_{W}$ is the channel number of the main stage. Source refers to the type of ViT from which the tracking model is adapted. The experiments are conducted on ITB [83].
@@ 5.4 Ablation Study and Analysis
1 17 320 PVT [15] Hier. 23M 42 FPS 57.5 %
2 23 320 PVT [15] Hier. 35M 24 FPS 58.2 %
3 12 768 ViT [13] plain 85M 102 FPS 61.1 %
4 24 1024 ViT [13] plain 305M 22 FPS 62.4 %
Table VII: Improvements of SuperSBT. ✔denotes adopting hierarchical structure, Masked Image Modeling (MIM) pre-training, and temporal modeling scheme.
  No. Hierarchical Pre-train Temporal. GOT-10k LaSOT
@@
  1 62.1 % 70.3% 55.7% 61.5%
2 65.9% 75.1% 58.2% 64.6%
3 71.0% 81.4% 66.5% 71.0%
4 64.2% 71.5 % 57.1 % 62.4 %
5 74.0 % 83.5 % 69.8% 75.6%
@@
Target-aware feature embedding. We explore the features of three different settings in two folds: maintaining spatial location information and identify the target from distractor objects. To begin with, we train three models that have only Cls head for localizing the target: (a) Correlation-embedded tracker. (b) Siamese correlation with SBT. (c) Siamese correlation with ResNet-50. We randomly jitter the search image around the target in the five hardest videos from the OTB [80] benchmark. We evaluate only the Cls map for localization.In Fig. 9, we observe that the true positive rate of the target ground truth demonstrates that (a) and (b) can preserve more spatial information compared to (c) CNN. The T-SNE/Cls map also exhibits the target-dependent characteristic of (a) features. The average negative objects (largest connected components) of (a) are higher than (b), indicating the effectiveness of the correlation embedding.
@@
Effect of architectural variants We investigate the effects of three architectural factors on the performance and inference speed of our method, namely the number of transformer blocks $N_{D}$ (depth), the channel dimension in each block $N_{W}$ (width), and the model structure (isotropic/hierarchical). Tab. VI shows that model performance depends very weakly on its shape (depth and width) but is significantly influenced by its scale (Models No.3/No.4 exhibit notably better performance than small-scale models). The observation that model No.1 has a smaller scale but lower speed highlights that the speed heavily hinges on the model shape.
Improvements of SuperSBT. In Tab. VII, we study the effects of three modifications made on SuperSBT, i.e., hierarchical structure, masked image modeling pre-training, and temporal modeling scheme. We observe that SuperSBT significantly outperforms the baseline, confirming the efficacy of these three modifications. Moreover, consistent performance gains can be obtained when each of the three modifications is applied individually, with the largest improvement yielded by masked image modeling pre-training.
6 Conclusion
In this work, we propose a novel fully transformer-based Single-Branch Tracking framework (SBT). Our SBT greatly simplifies the popular Siamese tracking pipeline by unifying feature extraction and correlation steps as one stage. Moreover, we conduct a systematic study on SBT tracking and summarize a bunch of effective design principles. Based on the summarized principles, we have developed an improved SuperSBT model for tracking. Enhanced by reasonable architecture variant design, masked image modeling pre-training, and temporal modeling, our SuperSBT model delivers superior results while raising the running speed with an even larger model capacity. Experiments on eight VOT benchmarks verify that Our SBT and SuperSBT variants achieve SOTA results while maintaining a simple architecture, showing a great potential to serve as a strong baseline tracker.

### v2

@@ IV Improved Single-Branch Tracking
IV-A Overview of Improvements
The overall architecture of SuperSBT is illustrated in Fig. 5. Following Principle A and Principle D, SuperSBT adopts a hierarchical transformer architecture (see Sec. IV-B) and independent layer design for shallow and deep model stages. Specifically, we stack local modeling layers (see Sec. IV-C) and patch merging layers in the first two stages, gradually reducing the number of tokens and expanding the channel dimension. It is worth noting that the initial two-stage processes involving the template and search image tokens are separated. Based on Principle B, we propose a unified relation modeling layer (see Sec. IV-D) for joint feature extraction and correlation in the third stage. Following Principle C and Principle E, we select the optimal architecture variants (see Sec. IV-I), such as layer number and channel dimension, to improve both performance and speed. Additionally, we chose patch-merging and a convolutional layer as our PaE method based on Principle F. In the final, SuperSBT is further enhanced by masked image modeling pre-training (see Sec. IV-F), a simple temporal modeling scheme (see Sec. IV-G), and an enhanced prediction head (see Sec. IV-H).
IV-B Hierarchical Architecture
@@
Figure 8: (a) The dynamic template mechanism in Stark [44] tracker. (b) Our temporal modeling scheme in SBT tracking.
IV-C Local Modeling Layer
Here, we present the local modeling layer, for the shallow model stage. The structure of the local modeling layer is illustrated in Fig. 6(a). Unlike the standard transformer layer[40, 14, 15], the Multi-head attention layer is replaced by the multi-layer perceptron to enrich the local representations in the shallow model stage. The remaining part of the local modeling layer follows the standard transformer layer. The data processing can be formulated as follows:
$\mathbf{f}_{att}&=\mathbf{f}+\text{MLP}(\text{LayerNorm}(\mathbf{f}),\quad\mathbf{f})\in\{z,x\}\\ \mathbf{f}_{out}&=\mathbf{f}_{att}+\text{MLP}(\text{LayerNorm}(\mathbf{f}_{att})),\quad\mathbf{f}\in\{z,x\}\end{aligned}$ (10)
@@ IV-C Local Modeling Layer
IV-D Unified Relation Modeling
@@
$\mathbf{f}_{zx}&=\text{Concat}(\mathbf{f}_{z},\mathbf{f}_{x}),\\ \mathbf{f}_{zx-att}&=\mathbf{f}_{zx}+\text{MHSA}(\text{LayerNorm}(\mathbf{f}_{zx})),\\ \mathbf{f}_{zx-out}&=\mathbf{f}_{zx-att}+\text{MLP}(\text{LayerNorm}(\mathbf{f}_{zx-att})),\end{aligned}$ (11)
@@
Based on Principle B, we do not adopt any specialized local attention (e.g., window attention), as it is harmful to parallel computing and inference speed. Moreover, the URM method can naturally learn the layer pattern of intra-/inter-image modeling for the whole network, which abandons the handcrafted placement of the FRM-SA/-CA layers. The information flow in single URM layer $\{\overrightarrow{zx},\overrightarrow{xz},\overrightarrow{zz},\overrightarrow{xx}\}$ is more comprehensive than that in single FRM-SA/-CA layer $\{\overrightarrow{zz},\overrightarrow{xx}\}$/$\{\overrightarrow{zx},\overrightarrow{xz}\}$. The computation Flops of attention are calculated as $4C^{2}L+2CL^{2}$, where $L$ is the number of feature tokens. The typical hyper-parameters are $\{64,256\}$ for token numbers $\{L_{z},L_{x}\}$, and $C=512$ for channel dimensions in the main stage. Thus, compared to the FRM layer ($0.34G$), the increased computation overhead of one URM layer ($0.55G$) is around $0.21$G, which has a negligible impact on the tracking efficiency.
IV-E Relative Position Encoding
@@ 3 $8\&16$
IV-F Masked Image Modeling Pre-training
@@
IV-G Temporal Modeling Scheme
We demonstrate our temporal modeling scheme to capture the target appearance changes over time. Similar to the dynamic template mechanism in Stark [44], we introduce a dynamically updated template sampled from intermediate frames as an additional input, as shown in Fig. 8(b). In the main stage of SueprSBT, triplet image inputs (first frame, dynamic template, and search frame) are concatenated and then sent to the FRM layers. The dynamic template is initialized from the first frame and updated by a predefined confidence threshold and time interval. More details are in the supplementary materials.
Our SuperSBT can directly update the image input while Stark updates the image features extracted from ResNet [12] backbone, and only the decoder part models the temporal variations. Thus, our SuperSBT learns the spatiotemporal relationship among templates and search regions more thoroughly than Stark.
IV-H Mix-MLP Prediction Head
In addition to the Convoluational-based head, we propose a Mix-MLP head, which can jointly model the dependency between the spatial and channel dimensions of the input features. The detailed analysis is in supplementary materials.
@@ IV-H Mix-MLP Prediction Head
IV-I Network Variants
Following the guidelines from Sec. III-D, five variants of SBT are described in Tab. II. It includes two base SBT variants and three SuperSBT variants. We place most layers in the third model stage based on Principle C. Moreover, we adjust the channel numbers and layer dimensions to maintain high speed and model performance based on Principle E.
IV-J Training Loss
@@
Method Year LaSOT [62] TrackingNet [19] GOT-10k [1] LaSOT-Ext [63]
AUC PNorm P AUC PNorm P AO SR0.5 SR0.75 AUC P
SuperSBT-Light ours 65.8 75.3 70.6 81.4 86.2 79.3 69.4 79.4 64.1 44.9 50.1
SuperSBT-Small ours 67.5 77.1 73.1 82.7 87.3 81.2 71.6 81.4 68.3 45.2 50.3
SuperSBT-Base ours 70.0 79.8 76.1 84.0 88.4 83.2 74.4 83.9 71.3 48.1 54.2
SuperSBT-Base-384 ours 72.8 82.5 78.6 84.8 88.9 83.7 75.5 84.3 72.4 50.7 57.9
Plain-SBT ours 65.7 75.4 70.1 81.2 86.1 80.3 69.2 79.1 64.0 44.2 49.3
Hi-SBT ours 65.3 74.7 70.3 81.0 85.6 79.0 69.5 79.8 63.9 44.6 49.6
SeqTrack [64] 2023 69.9 79.7 76.3 83.3 88.3 82.2 74.7 84.7 71.8 49.5 56.3
ROMTrack [65] 2023 69.3 78.8 75.6 83.6 88.4 82.7 72.9 82.9 70.2 48.9 55.0
GRM [66] 2023 69.9 79.3 75.8 84.0 88.7 83.3 73.4 82.9 70.4 - -
OStrack [47] 2022 69.1 78.7 75.2 83.1 87.8 82.0 71.0 80.4 68.2 47.4 53.3
SimTrack [49] 2022 69.3 78.5 - 82.3 86.5 - 68.6 78.9 62.4 - -
Mixformer [48] 2022 69.2 78.7 74.7 83.1 88.1 81.6 70.7 80.0 67.8 - -
SwinTrack [45] 2022 67.2 - 70.8 81.1 - 78.4 71.3 81.9 64.5 49.1 55.6
TransT [4] 2022 64.9 73.8 69.0 81.4 86.7 80.3 67.1 76.8 60.9 - -
ToMP50 [25] 2022 67.6 - - 81.2 86.4 78.9 - - - 45.9 -
CSWinTT [67] 2022 66.2 75.2 70.9 81.9 86.7 79.5 69.4 78.9 65.4 - -
UTT [68] 2022 64.6 - 67.2 79.7 - 77.0 67.2 76.3 60.5 - -
ARDiMPsuper [56] 2021 65.3 73.2 68.0 80.5 85.6 78.3 70.1 80.0 64.2 - -
TrDiMP [43] 2021 63.9 - 61.4 78.4 83.3 73.1 68.8 80.5 59.7 - -
TrSiam [43] 2021 62.4 - 60.0 78.1 82.9 72.7 67.3 78.7 58.6 - -
STMTrack [33] 2021 60.6 69.3 63.3 80.3 85.1 76.7 64.2 73.7 57.5 - -
SiamBAN-ACM [29] 2021 57.2 - - 75.3 81.0 71.2 - - - - -
SiamGAT [69] 2021 53.9 63.3 53.0 - - - 62.7 74.3 48.8 - -
DSTrpn [70] 2021 43.4 51.3 - 64.9 58.9 - - - - - -
SiamR-CNN [71] 2020 64.8 72.2 - 81.2 85.4 80.0 64.9 72.8 59.7 - -
Ocean [72] 2020 56.0 65.1 56.6 - - - 61.1 72.1 47.3 - -
KYS [73] 2020 55.4 63.3 - 74.0 80.0 68.8 63.6 75.1 51.5 - -
DCFST [24] 2020 - - - 75.2 80.9 70.0 63.8 75.3 49.8 - -
SiamFC++ [27] 2020 54.4 62.3 54.7 75.4 80.0 70.5 59.5 69.5 47.9 - -
PrDiMP [74] 2020 59.8 68.8 60.8 75.8 81.6 70.4 63.4 73.8 54.3 - -
SiamAttn [32] 2020 56.0 64.8 - 75.2 81.7 - - - - - -
MAML [36] 2020 52.3 - - 75.7 82.2 72.5 - - - - -
D3S [75] 2020 - - - 72.8 76.8 66.4 59.7 67.6 46.2 - -
SiamCAR [28] 2020 50.7 60.0 51.0 - - - 56.9 67.0 41.5 - -
SiamBAN [29] 2020 51.4 59.8 52.1 - - - - - - - -
DiMP [9] 2019 56.9 65.0 56.7 74.0 80.1 68.7 61.1 71.7 49.2 39.2 45.1
SiamPRN++ [7] 2019 49.6 56.9 49.1 73.3 80.0 69.4 51.7 61.6 32.5 34.0 39.6
ATOM [3] 2019 51.5 57.6 50.5 70.3 77.1 64.8 55.6 63.4 40.2 37.6 43.0
ECO [22] 2017 32.4 33.8 30.1 55.4 61.8 49.2 31.6 11.1
MDNet [76] 2016 39.7 46.0 37.3 60.6 70.5 56.5 29.9 30.3 9.9 27.9 31.8
SiamFC [6] 2016 33.6 42.0 33.9 57.1 66.3 53.3 34.8 35.3 9.8 - -
@@
V Experiments
@@ V-A Implementation Details
V-B Evaluation on TrackingNet, LaSOT, and GOT-10k
@@
TrackingNet. TrackingNet [19] is a large-scale tracking dataset that covers diverse object classes and scenes. Its test set contains 511 sequences of publicly available ground truth. We submit our tracker’s outputs to the official online evaluation server and report the Success (AUC) and Precision (P and PNorm) results in Tab. III. As can be seen, SuperSBT-Base-384 achieves the best performance, achieving 84.0%, 88.4%, and 83.2% in terms of AUC, $\rm{P}_{Norms}$, and P, respectively. Besides, in terms of AUC, SuperSBT-Base surpasses the fully-transformer tracker OStrack [47] by 0.9% and two transformer-based trackers, TrDiMP and TrSiam, by 5.6% and 5.9%, respectively.
GOT-10k. The GOT-10k [1] dataset contains 10k sequences for training and 180 for testing. We follow the defined protocol presented in [1] and submit the tracking outputs to the official evaluation server. The results (i.e., AO and SRT) are reported in Tab. III. It can be seen that our SuperSBT-Base-384 achieves the best performance. SuperSBT-Base and SuperSBT-Light obtain 74.4% and 69.4% AO, respectively, leading previous methods by significant margins. SuperSBT-Base achieves comparable performance in terms of AO at a much faster speed compared to SeqTrack [64] under the same input resolution (81 FPS vs 37 FPS). Furthermore, SuperSBT-Base outperforms the recent transformer tracker GRM [66] by 1.0% in terms of AO while running at a faster speed (81 FPS vs 40 FPS).
LaSOT-Ext. LaSOT-Ext [63] is a recently released dataset with 150 video sequences and 15 object classes, which contains highly challenging scenarios. Our SuperSBT-Base still outperforms the SOTA trackers OStrack and achieves the top rank with $50.7\%$ AO with the $384$ input setting. Our small variant achieves comparable performance to ToMP50 ($45.2\%$ vs. $45.9\%$) while running much faster (110 FPS vs. 30 FPS). It validates the effectiveness and efficiency of our method.
@@
Method LaSOT [62] GOT-10k [1] NFS [80] Speed in GPU
AUC PNorm P AO SR0.5 SR0.75 AUC FPS
SuperSBT-Light 65.8 75.3 70.6 69.4 79.4 64.1 65.1 141
SuperSBT-Small 67.5 77.1 73.1 71.6 81.4 68.3 66.2 110
HiT-Tiny [83] 54.8 60.5 52.9 52.6 59.3 42.7 53.2 204
HiT-Small [83] 60.5 68.3 61.5 62.6 71.2 54.4 61.8 192
HiT-Base [83] 64.6 73.3 68.1 64.0 72.1 58.1 63.6 175
E.T.Track [84] 59.1 - - - - - 59.0 -
FEAR-XS [85] 53.5 - 54.5 61.9 - - 61.4 -
FEAR-M [85] 54.6 - 55.6 62.3 - - 62.2 -
LightTrack-Mobile [86] 53.8 - 53.7 61.1 71.0 - - -
LightTrack-LargeA [86] 55.0 - 55.2 61.5 72.3 - - -
Table VI: State-of-the-art comparison with lightweight trackers on LaSOT, GOT-10k, and NFS benchmarks. The best three are shown in red, blue, and green fonts.
V-C Evaluation on Other Datasets
@@
V-D Comparing to Lightweight Trackers
As shown in Tab. VI, compared to the specifically designed lightweight trackers (HiT [83], FEAR [85], E.T.Track [84], and LightTrack [86]), our SuperSBT-light/small tracker outperforms them by a great margin in three benchmarks, while still having comparable speed (over 100 FPS). This indicates that our lightweight variants can achieve marvelous performance as well as maintain competitive efficiency.
V-E Ablation Study and Analysis
@@
1 17 320 PVT [15] Hier. 23M 42 FPS 57.5
2 23 320 PVT [15] Hier. 35M 24 FPS 58.2
3 12 768 ViT [13] plain 85M 102 FPS 61.1
4 24 1024 ViT [13] plain 305M 22 FPS 62.4
Table VII: Ablations on the effects of architecture variants: network depth $N_{D}$, width $N_{W}$, and model structure (plain/hierarchical). $N_{W}$ is the channel number of the main stage. Source refers to the type of ViT from which the tracking model is adapted. The experiments are conducted on ITB [88].
  No. Hierarchical Pre-train Temporal. MixMLP. GOT-10k LaSOT
@@
  1 62.1 70.3 55.7 61.5
2 65.9 75.1 58.2 64.6
3 71.0 81.4 66.5 71.0
4 64.2 71.5 57.1 62.4
5 63.5 71.8 57.8 63.9
6 74.2 83.6 69.9 76.0
Table VIII: Improvements of SuperSBT. ✔denotes adopting hierarchical structure, masked image modeling pre-training, temporal modeling, and Mix-MLP head.
@@
Target-aware feature embedding. We explore the features of three different settings in two folds: maintaining spatial location information and identifying the target from distractor objects. To begin with, we train three models that have only Cls head for localizing the target: (a) Correlation-embedded tracker. (b) Siamese correlation with SBT. (c) Siamese correlation with ResNet-50. We randomly jitter the search image around the target in the five hardest videos from the OTB [81] benchmark. We evaluate only the Cls map for localization.In Fig. 10, we observe that the true positive rate of the target ground truth demonstrates that (a) and (b) can preserve more spatial information compared to (c) CNN. The T-SNE/Cls map also exhibits the target-dependent characteristic of (a) features. The average negative objects (largest connected components) of (a) are higher than (b), indicating the effectiveness of the correlation embedding.
@@
Effect of architectural variants. We investigate the effects of three architectural factors on the performance and inference speed of our method, namely the number of transformer blocks $N_{D}$ (depth), the channel dimension in each block $N_{W}$ (width), and the model structure (isotropic/hierarchical). Tab. VII shows that model performance depends very weakly on its shape (depth and width) but is significantly influenced by its scale (Models No.3/No.4 exhibit notably better performance than small-scale models). The observation that model No.1 has a smaller scale but lower speed highlights that the speed heavily hinges on the model shape.
Improvements of SuperSBT. In Tab. VIII, we study the effects of three modifications made to SuperSBT: hierarchical structure, masked image pre-training, Mix-MLP head, and temporal modeling scheme. We observe that SuperSBT significantly outperforms the baseline, confirming the efficacy of these three modifications. Moreover, consistent performance gains can be obtained when each of the five modifications is applied individually, with the largest improvement yielded by masked image pre-training.
VI Conclusion
In this work, we propose a novel fully transformer-based Single-Branch Tracking framework (SBT). Our SBT greatly simplifies the popular Siamese tracking pipeline by unifying feature extraction and correlation steps as one stage. Moreover, we conduct a systematic study on SBT tracking and summarize a bunch of effective design principles. Based on the summarized principles, we have developed an improved SuperSBT model for tracking. Enhanced by reasonable architecture variant design, masked image modeling pre-training, and temporal modeling, our SuperSBT model delivers superior results while raising the running speed with an even larger model capacity. Experiments on eight VOT benchmarks verify that our SBT and SuperSBT variants achieve SOTA results while maintaining a simple architecture, showing great potential to serve as a strong baseline tracker.